Calculate the present value of $9,000 received five years from today if your investments pay (Do not round intermediate calculations. Round your answers to 2 decimal places. (e.g., 32.16))


           Present Value
a. 6 percent compounded annually $     
b. 8 percent compounded annually     
c. 10 percent compounded annually     
d. 10 percent compounded semiannually     
e. 10 percent compounded quarterly
Answer #1

Present value = Future value x Present value factor

Present value factor

= 1 / (1 + r) ^ n

Where,

r = Rate of interest

n = Number of periods

When interest is compounded semi-annually or quarterly, interest rate is divided by 2 and 4 respectively and time period is multiplied by 2 and 4 respectively

a)

Present value factor

= 1 / 1.06 ^ 5

= 1 /1.338225

= 0.747258

So, Present value = $9,000 x 0.747258

= $6,725.32

b)

Similarly,

Present value factor

= 1 / 1.08^5

= 1 / 1.469328

= 0.680583

So, Present value = $9,000 x 0.680583

= $6,125.25

c)

Present value factor

= 1 / 1.10^5

= 1 / 1.61051

= 0.620921

So, Present value = $9,000 x 0.620921

= $5,588.29

d)

Since interest is compounded semi-annually, interest rate will be 10 / 2 = 5% per period and time period will be 5 x 2 = 10 semi-annual periods

So, Present value factor

= 1 / 1.05^10

= 1 / 1.628895

= 0.613913

So, Present value = $9,000 x 0.613913

= $5,525.22

e)

Since interest is compounded quarterly, interest rate will be 10 / 4 = 2.5% per period and time period will be 5 x 4 = 20 quarters

So, Present value factor

= 1 / 1.025^20

= 1 /1.638616

= 0.610271

So, Present value = $9,000 x 0.610271

= $ 5,492.44